C6H6 + Cl2 ---------> C6H5Cl + HCl
What mass of HCl will be produced, if 2.5 g of C6H6 reacts with 1.5 g of C6H5Cl?
1
Expert's answer
2020-11-24T05:54:19-0500

Solution.

n(C6H6) = 0.032 mol

n(C6H6) = n(C6H5Cl) = 0.032 mol

m(C6H5Cl) = 3.6 g

"\phi = \frac{1.5}{3.6} \times 100 \% = 41.67 \%"

m(HCl) = 1.168 g

m(HCl)prac = 0.49 g

Answer:

m(HCl)prac = 0.49 g